# Beyond Image Prior: Embedding Noise Prior into Conditional Denoising Transformer

### v1

@@ Abstract
Index Terms:
Image denoising, vision Transformer, noise modeling, conditional optimization.
@@ 1 Introduction
In general, natural images embody strong priors for visual perception [1, 2], such as repetitive textures and continuous edges, which are more readily inferred than the seemingly random presence of noise. Thus, a prevailing strategy among many existing learning-based denoising methods [3, 4, 5, 6, 7, 8, 9] involves developing a unified model capable of generalizing from a vast collection of noisy-clean pairs. This process typically formulates optimization as:
@@
Figure 1: Illustration of denoising model optimization: (a) unconditional optimization space with image prior [8, 9]; (b) conditional optimization with noise prior and image prior in this work.
@@
1) The difficulty and cost associated with gathering large-scale noisy-clean image datasets. Learning sophisticated image priors necessitates a model with substantial capacity, which is hindered by the challenges of acquiring clean images. Clean image acquisition typically requires long exposures in static scenes [10, 11, 12] or involves complex alignment procedures [10].
@@
To address these issues, we propose segmenting the singular unconditional optimization space into multiple subspaces by incorporating reliable noise priors alongside the complicated image priors, as illustrated in Fig. 1b.This approach recognizes that a noisy observation is influenced both by the scene (image prior) and the imaging environment (noise prior), making it logical to infer the noise prior for optimizing the denoising process. Consequently, the optimization space for a conditional denoising model should comprise various independent and complete subspaces, each conditioned on specific priors, and can be represented as:
@@
Distinct from traditional models that rely on given conditional embeddings (e.g., cGAN [13], cVAE [14] and DnCNN [15]) or implicit noise prediction (e.g., VDN [16] and CVF-SID [17]), a conditional denoising model must adeptly estimate an explicit noise prior from a single noisy observation and distinctly address the noise and image priors based on their independence. In essence, as depicted in Fig. 1b, the principle of a conditional denoising model lies in its ability to navigate the generation of pixels by harnessing implicit natural image priors to shape the optimization landscape, while also leveraging explicit sensor noise priors to precisely target the optimization’s focus.
@@
  • By exploring the noise statistics concealing in the latent space, we propose a novel Condformer that leverages the estimated noise prior within a conditional self-attention module. This design represents a pioneering effort to incorporate prior knowledge into a Transformer-like architecture for denoising.
@@ 2.1 Noise Modeling
In general, the signal-independent Gaussian distribution is regarded as a theoretically common hypothesis of noise modeling, and has derived numerous supervised image denoising methods to handle the widely-used additive white Gaussian noise (AWGN). However, the real noise model is more sophisticated. Particularly, due to the characteristics of imaging sensor, the practical noise could be explicitly modeled on raw sensors, and the corresponding raw sensor noise commonly consists of the signal-dependent shot noise and the signal-independent read noise. Typically, the Poisson-Gaussian noise model [18] was employed to characterize the distribution of this raw sensor noise and has inspired numerous advances in realistic noise synthesis and real image denoising [19, 20, 21, 17]. Beyond the hypothesis on building the distribution of noise, to match the noise model in more complex imaging environment on various devices, multi-frame calibration [22, 18, 21] and variance-stabilizing transformations [23, 24] technologies were presented to refine the noise model for different sensors. Furthermore, some works attempt to employ the calibrated noise parameter of sensor to explicitly synthesize noisy samples [21, 25, 26], or guide the network optimization [16, 17, 27, 28] for training a denoising model. Instead of explicitly modeling noise from a certain distribution, learnble noise modeling methods recently have been raised with the development of generative models, such as variational bayes [29], generative adversarial networks [30, 31], and normalizing flows [32, 33, 34].
@@ 2.2 Image Denoising
After decades of development, image denoising methods are generally divided into model-based and learning-based. Model-based denoising methods aim to model the characteristics of natural images as a regularization prior to iteratively optimize a well-designed model. The representative regularization priors include total variation [35], sparsity [36], non-local self-similarity [37, 38], and external statistical priors [39]. On the other hands, learning-based methods attempt to reconstruct a clean image from the noisy observation with an end-to-end learnable model, which is trained from large-scale noisy-clean pairs. Initially, with the development of convolutional neural networks (CNN), CNN-based denoising methods [15, 40, 3, 41, 5, 4] have received significant advances in learning an end-to-end mapping from the noisy observations to clean targets. In essence, due to the nature of CNN in local visual perception, the key of these CNN-based denoising methods is to learn how a pixel is generated from a corrupted one and its neighbors in local perception region, namely, to learn the image prior in local perceptions, such as details in texture, smoothness in flat. Nevertheless, some contextual image priors are difficult to capture in local perceptions, e.g., objects, structural information, edges and repeated textures. Recently, considering these image priors in non-local or global visual perception, several Transformer-based [42, 6, 7], MLP-based [43, 44, 45] and Mamba-based [46, 47] denoising methods have attracted increasing attentions and achieved remarkably superior performances against other existing CNN-based methods. In particular, to capture long-range feature dependencies, pyramid [48], rectangle-window[49], chaotic-window [50], sparse [51] and anchored-stripe [52] self-attention mechanisms have been explored. Yet capturing spatial correspondence commonly causes quadratically increasing computational loads as the resolution increases, then efficient Transformer-based denoisers recently achieve growing concerns. Specifically, hierarchical U-shape architecture [8] and channel-wise self-attention mechanism [9] were proposed to reduce the unbearable computational loads from increasing spatial resolutions.
@@
Except for the evolution of denoiser architectures, a blind denoising strategy with stronger generalization is essential for practical applications due to the unknowability and the diversity of noise in real scene. In the type of aforementioned supervised learning-based methods, numerous noisy-clean pairs with various noise levels are employed to train an unified denoiser [15, 53, 9, 54]. Besides, to adapt for real scenes with only noisy observations, self-supervised denoising methods [2, 17, 55] have been raised by learning implicit representation from image priors. However, they often fall short in scenarios requiring high accuracy, complex noise handling, and stable convergence.
@@ 2.3 Vision Transformer
The self-attention mechanism [42] in Transformers facilitates learning long-range dependencies, leading to significant success in natural language processing [42, 56] and computer vision tasks [57, 58]. For high-level vision, ViT [57] demonstrated the Transformer’s effectiveness in non-local visual perception and high-accuracy object recognition, sparking the development of more effective and efficient Transformer architectures for various vision tasks like object detection [58, 59], semantic segmentation [60, 61], and low-level vision [6, 9].
Especially in low-level vision, self-attention mechanism was firstly utilized to transfer relevant textures in reference-based super-resolution [62]. More generally, IPT [6] later introduced a multi-task image processing model using standard Transformer architecture with tokenized inputs. By integrating the advantage of local attention mechanism of CNN and long-range dependency of Transformer, Swin Transformer [58] was proposed by introducing the shifted window scheme and was applied into image restoration tasks [7]. To alleviate the limitation of Swin Transformer in receptive fields, cross aggregation Transformer (CAT) [49] was proposed by aggregating features cross different windows to expand the receptive field. Despite their effectiveness, these Transformers are computationally intensive as calculating the spacial cross-covariance of large-scale tokens. To address this, Uformer [8] was proposed by building a hierarchical U-shape architecture with locally-enhanced window Transformer blocks. Besides, by calculating the channel-wise cross-covariance, an efficient Restormer [9] was proposed and achieved state-of-the-art performances in several image restoration tasks.
Nonetheless, these existing Transformers rely on large-scale datasets with perfect labels and unconditional optimization, which is commonly redundant. Instead, inspired by the conditional text generation with Transformer [63, 64], we aim to explore a conditional Transformer for image denoising.
Figure 2: Investigation on independence of the image prior and the noise prior. We select three noisy-clean pairs from SIDD training dataset, which are captured with different image priors (related to scene and brightness) or noise priors (related to ISO level and shutter speed). By investigating the local variance of noisy residues and the local mean of clean image, the statistical results show that the noise prior depends on only the camera settings, but little on the image prior.
@@ 3.1.1 Noise Formation Model
Although the noise in a processed sRGB image is generally complex to explicitly analyze due to the nonlinearity of image signal processing (ISP), the raw noise formation model of a digital sensor in camera is well understood [65]. In particular, the raw noise in a camera primarily consists of the shot noise during photon-to-electron conversion and the read noise during electron-to-digital conversion [25].
@@
This Poisson-Gaussian can be further treated as a heteroscedastic Gaussian distribution [18], and the final raw digital sensor signals can be formulated as
@@ 3.1.2 Independence of Noise Prior and Image Prior
However, it is infeasible to calculate the pixel-wise variances on a single image. We randomly sample $1000\times 16\times 16$ local raw noisy-clean patch pairs, and calculate the local variances $\sigma^{2}$ of the noisy residues and the local means $\mu$ of the clean patches, to further approximate the corresponding pixel-wise statistic of noise. As formulated in Eq.(5), the pixel-wise variance $\sigma^{2}$ of noises should be proportional to the intensity $\mu$ of illuminances. Consequently, as the statistical results shown in Fig. 2, its slope and intercept represent $\sigma^{2}_{\text{s}}$ and $\sigma^{2}_{\text{r}}$, respectively.
@@
Figure 3: Pipeline of LoNPE algorithm. The dynamic range of the input raw noisy image has been normalized into $[0,1]$.
@@ 3.2 Locally Noise Prior Estimation
Specifically in practical usage, we can only capture a corrupted noisy observation and commonly have little additional information on the environmental illumination, sensor technology, and photography settings. Therefore, the independence of noise prior and image prior is a critical principle that we can use to estimate the noise prior from a single noisy image.
@@
where $\mathcal{I}(\cdot)$ indicates image prior model implemented as an onefold denoising model, $\mathcal{O}(i)$ denotes the local neighbors of location $i$.
@@
in other words, the image prior is negligible in a local smooth patch, and is represented as the mean value of pixels in local region. Thus, on the local neighbor locations $\mathcal{O}(i)$, the statistical distribution of $\{\mathbf{I}_{i}\}_{i\in\mathcal{O}(i)}$ should be approximately same, and could be utilized to estimate the noise prior as its independency on the image prior.
As illustrated in Fig. 3, we firstly preprocess the image to restrict its theoretical value range into $[0,1]$, then partition it into a group of local patches, and select the smooth patches to eliminate the interference of image prior. Next, we calculate these patches’ statistical values of mean and variance, to finally estimate the noise prior ($\sigma_{\text{s}}$, $\sigma_{\text{r}}$) with a simple least square optimization solver.
@@ 3.2.1 LoNPE Algorithm
Due to the variety of sensor bit depth $B$ (e.g., $B=10$ in iPhone 7 camera, $B=14$ in Canon 80D camera), the theoretical value range of raw digital sensor image $[0,2^{B}]$ might be different. For generally analyzing, we preprocess the raw image by normalizing its theoretical value range into $[0,1]$, as
$\mathbf{I}=\mathbf{I}\cdot 2^{-B}$ (8)
@@
Figure 4: Framework of LoNPE network. The backbone of network is same as the one in DoTNet [66].
@@
[…]
@@ 3.2.2 LoNPE Network
Inspired by the concept of noise representation in [66], that the noise level could be represented as an explicit parameter and is easy to learn by an external neural network. We build a learnable CNN model to predict the noise prior $(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})\rightarrow(\sigma_{\text{s}},\sigma_{\text{r}})$ from a single sRGB noisy color image $\mathbf{I}_{\text{rgb}}$, which is formulated as
@@
where $\Phi_{\text{LoNPE}}(\cdot)$ denotes the noise prior estimation network, and called “LoNPE network”. The framework of LoNPE network is shown in Fig. 4, consisting of a backbone as the DoTNet in [66] for feature extraction and two fully-connected layers for decision. Due to the physical characteristics of shot noise (photon-to-electron conversion rate) and read noise (quantization range of digital signal), the output range of $\sigma_{\text{s}}$ and $\sigma_{\text{r}}$ are limited to $[0,1]$.

### v2

# Beyond Image Prior: Embedding Noise Prior into Latent Space of Conditional Denoising Transformer

@@ Abstract
keywords:
Image denoising, Vision Transformer, Noise modeling, Conditional optimization
@@ 1 Introduction
Figure 1: Illustration of denoising model optimization: (a) unconditional optimization space with image prior $x$ 67, 21; (b) conditional optimization with noise prior $z$ and image prior $x$ in this work.
In general, natural images embody strong priors for visual perception 51, 28, such as repetitive textures and continuous edges, which are more readily inferred than the seemingly random presence of noise. Thus, a prevailing strategy among many existing learning-based denoising methods 71, 72, 67, 42, 21 involves developing a unified model capable of generalizing from a vast collection of noisy-clean pairs. This process typically formulates optimization as:
@@
1) The difficulty and cost associated with gathering large-scale noisy-clean image datasets. Learning sophisticated image priors necessitates a model with substantial capacity, which is hindered by the challenges of acquiring clean images. Clean image acquisition typically requires long exposures in static scenes 1, 3, 46 or involves complex alignment procedures 1.
@@
To address these issues, we propose segmenting the singular unconditional optimization space into multiple subspaces by incorporating reliable noise priors alongside the complicated image priors. As illustrated in Fig. 1b, this approach recognizes that a noisy observation is influenced both by the scene (image prior) and the imaging environment (noise prior), making it logical to infer the noise prior for optimizing the denoising process. Consequently, the optimization space for a conditional denoising model should comprise various independent and complete subspaces, each conditioned on specific priors, and can be represented as:
@@
Distinct from traditional models that rely on given conditional embedding (e.g., FFDNet 72) or implicit noise prediction (e.g., VDN 63 and CVF-SID 43), a conditional denoising model must adaptively estimate an explicit noise prior from a single noisy observation and distinctly address the noise and image priors based on their independence, rather than concatenate the image and noise parameter directly as their mismatching is a critical obstacle for improving denoising performance. In essence, as depicted in Fig. 1b, the principle of a conditional denoising model lies in its ability to navigate the generation of pixels by harnessing implicit natural image priors to shape the optimization landscape, while also leveraging explicit sensor noise priors to precisely target the optimization’s focus.
@@
  • By exploring the noise statistics concealing in the latent space, we propose a novel Condformer that leverages the estimated noise prior within a conditional self-attention module. This design represents a pioneering effort to incorporate prior knowledge into a Transformer-like architecture for denoising, and alleviates the mismatching issue in existing conditional denoisers.
@@ 2.1 Noise Modeling
In general, the signal-independent Gaussian distribution is regarded as a theoretically common hypothesis of noise modeling, and has derived numerous supervised image denoising methods to handle the widely-used additive white Gaussian noise (AWGN). However, the real noise model is more sophisticated. Particularly, due to the characteristics of imaging sensor, the practical noise could be explicitly modeled on raw sensors, and the corresponding raw sensor noise commonly consists of the signal-dependent shot noise and the signal-independent read noise.
Typically, the Poisson-Gaussian noise model 18 was employed to characterize the distribution of this raw sensor noise and has inspired numerous advances in realistic noise synthesis and real image denoising 33, 22, 54. Beyond the hypothesis on building the distribution of noise, to match the noise model in more complex imaging environment on various devices, multi-frame calibration 54 and variance-stabilizing transformations 38, 39, 29 technologies were presented to refine the noise model for different sensors. Furthermore, some works attempt to employ the calibrated noise parameter of sensor to explicitly synthesize noisy samples 54, 57, 17, or guide the network optimization 43, 64 for training a denoising model. Instead of explicitly modeling noise from a certain distribution, learnble noise modeling methods recently have been raised with the development of generative models, such as variational bayes 77, generative adversarial networks 9, and normalizing flows 40.
@@ 2.2 Image Denoising
After decades of development, image denoising methods are generally divided into model-based and learning-based. Model-based denoising methods aim to model the characteristics of natural images as a regularization prior to iteratively optimize a well-designed model. The representative regularization priors include total variation 47, sparsity 58, non-local self-similarity 6, 15, external statistical priors 61, and Huber function 48.
On the other hands, learning-based methods attempt to reconstruct a clean image from the noisy observation with an end-to-end learnable model, which is trained from large-scale noisy-clean pairs.
2.2.1 Non-blind image denoising
@@
Initially, with the development of convolutional neural networks (CNN), CNN-based denoising methods 71, 72, 4, 42, 75, 66, 68, 44 have received significant advances in learning an end-to-end mapping from the noisy observations to clean targets. In essence, due to the nature of CNN in local visual perception, the key of these CNN-based denoising methods is to learn how a pixel is generated from a corrupted one and its neighbors in local perception region, namely, to learn the image prior in local perceptions, such as details in texture, smoothness in flat. Nevertheless, some contextual image priors are difficult to capture in local perceptions, e.g., objects, structural information, edges and repeated textures. Recently, considering these image priors in non-local or global visual perception, several Transformer-based 11, 32, MLP-based 49, 50 and Mamba-based 21 denoising methods have attracted increasing attentions and achieved remarkably superior performances against other existing CNN-based methods. In particular, to capture long-range feature dependencies, pyramid 42, rectangle-window 76, chaotic-window 60, sparse 70 and anchored-stripe 30 self-attention mechanisms have been explored. Yet capturing spatial correspondence commonly causes quadratically increasing computational loads as the resolution increases, then efficient Transformer-based denoisers recently achieve growing concerns. Specifically, hierarchical U-shape architecture 56 and channel-wise self-attention mechanism 67 were proposed to reduce the unbearable computational loads from increasing spatial resolutions.
2.2.2 Blind image denoising
@@
Except for the evolution of denoiser architectures, a blind denoising strategy with stronger generalization is essential for practical applications due to the unknowability and the diversity of noise in real scene. In the type of aforementioned supervised learning-based methods, numerous noisy-clean pairs with various noise levels are employed to train an unified denoiser71, 73, 67, 14. To improve the performance, conditional embedding like noise variance map is concatenated with the noisy image in the head of denoiser, to guide model handling a specific given 72 or predicted 64 noise level. However, mismatch of image and noise level is a critical obstacle of these methods for denoising performance.
Besides, to adapt for real scenes with only noisy observations, self-supervised denoising methods 28, 43 were raised by learning implicit representation from image priors. However, they often fall short in scenarios requiring high accuracy, complex noise handling, and stable convergence.
@@ 2.3 Vision Transformer
The self-attention mechanism 52 in Transformers facilitates learning long-range dependencies, leading to significant success in computer vision tasks 16, 34. ViT 16 demonstrated the Transformer’s effectiveness in non-local visual perception and high-accuracy object recognition, sparking the development of more effective and efficient Transformer architectures for various vision tasks like object detection 34, 23, semantic segmentation 69, 74, and low-level vision 11, 67, 42.
Especially in low-level vision, self-attention mechanism was firstly utilized to transfer relevant textures in reference-based super-resolution 62. More generally, IPT 11 later introduced a multi-task image processing model using standard Transformer architecture with tokenized inputs. By integrating the advantage of local attention mechanism of CNN and long-range dependency of Transformer, Swin Transformer 34 was proposed by introducing the shifted window scheme and was applied into image restoration tasks 32. To alleviate the limitation of Swin Transformer in receptive fields, cross aggregation Transformer (CAT) 76 was proposed by aggregating features cross different windows to expand the receptive field. In addition, attention retractable Transformer (ART) 70 was presented to capture local and global receptive field simultaneously. GRL 30 was proposed to explicitly model image hierarchies in global, regional, and local range dependencies. Despite their effectiveness, these Transformers are computationally intensive as calculating the spacial cross-covariance of large-scale tokens. To address this, Uformer 56 was proposed by building a hierarchical U-shape architecture with locally-enhanced window Transformer blocks. Besides, by calculating the channel-wise cross-covariance, an efficient Restormer 67 was proposed and achieved state-of-the-art performances in several image restoration tasks.
Nonetheless, these existing Transformers rely on large-scale datasets with perfect labels and unconditional optimization, which is commonly redundant. Instead, inspired by the conditional text generation with Transformer 24, 78, we aim to explore a conditional Transformer for image denoising.
@@ 3.1.1 Noise Formation Model
Although the noise in a processed sRGB image is generally complex to explicitly analyze due to the nonlinearity of image signal processing (ISP), the raw noise formation model of a digital sensor in camera is well understood 5. In particular, the raw noise in a camera primarily consists of the shot noise during photon-to-electron conversion and the read noise during electron-to-digital conversion 57.
Figure 2: Investigation on independence of the image prior and the noise prior. We select 8 noisy-clean pairs from SIDD training dataset, which are captured with different image prior (related to scene and brightness) or noise prior (related to ISO level and shutter speed). By investigating the local variance of noisy residues and the local mean of clean image, the statistical results show that the noise prior depends on only the camera settings, but little on the image prior.
@@
This Poisson-Gaussian can be further treated as a heteroscedastic Gaussian distribution 18, and the final raw digital sensor signals can be formulated as
@@ 3.1.2 Independence of Noise Prior and Image Prior
However, it is infeasible to calculate the pixel-wise variances on a single image. We randomly sample 1000×16×161000$\times$ 16$\times$ 161000 × 16 × 16 local raw noisy-clean patch pairs, and calculate the local variances $\sigma^{2}$ of the noisy residues and the local means $\mu$ of the clean patches, to further approximate the corresponding pixel-wise statistic of noise. As formulated in Eq.(5), the pixel-wise variance $\sigma^{2}$ of noises should be proportional to the intensity $\mu$ of illuminances. Consequently, as the statistical results shown in Fig. 2, its slope and intercept represent $\sigma^{2}_{\text{s}}$ and $\sigma^{2}_{\text{r}}$, respectively.
@@
Figure 3: Pipeline of the LoNPE algorithm. The dynamic range of raw noisy image is normalized into [0,1][0,1][ 0 , 1 ].
@@ 3.2 Locally Noise Prior Estimation
Similar to conventional noise parameter estimation task, which has seen significant progress with methods evolving to address challenges in accuracy and robustness. Different from the Gaussian noise parameter estimation 10, 53, 27, 45, Poisson-Gaussian noise parameter estimation is more complicated as the noise variance is signal-dependent 18. Existing methods dedicate to solve the image variance of noisy observation via iterative variance-stabilization 39, maximum likelihood estimation 33, or learning-based CNN 7. But, challenges remain in handling real cases, especially within strong image prior such as rich textures.
Specifically in practical usage, as we can only capture a corrupted noisy observation and commonly have little additional information on the environmental illumination, sensor technology, and photography settings, noise prior is basically equivalent to the above noise parameter. To tackle the above challenge, the independence of noise prior and image prior is a critical principle that we can use to estimate the noise prior from a single noisy image.
@@
where $\Phi(\cdot)$ indicates image prior model implemented as an onefold denoising model, $\mathcal{O}(i)$ denotes the local neighbors of location $i$.
@@
as shown in Fig. 3, smooth patches exhibit more concentrated histogram distributions, allowing for a more precise characterization of the overall intensity within each patch. This indicates that the image prior is negligible in a local smooth patch and can be effectively represented as the mean value of the pixels within the local region. Thus, on the local neighbor locations $\mathcal{O}(i)$, the statistical distribution of $\{\mathbf{I}_{i}\}_{i\in\mathcal{O}(i)}$ should be approximately same, and could be utilized to estimate the noise prior as its independency on the image prior.
In particular, we firstly preprocess the image to restrict its theoretical value range into [0,1][0,1][ 0 , 1 ], then partition it into a group of local patches, and select the smooth patches to eliminate the interference of image prior. Next, we calculate these patches’ statistical values of mean and variance, to finally estimate the noise prior ($\sigma_{\text{s}}$, $\sigma_{\text{r}}$) with a simple least square optimization solver.
@@ 3.2.1 LoNPE Algorithm
Due to the variety of sensor bit depth $B$ (e.g., $B=10$ in iPhone 7 camera, $B=14$ in Canon 80D camera), the value range of raw digital sensor image $[0,2^{B}]$ might be different. For generally analyzing, we preprocess the raw image by normalizing its theoretical value range into [0,1][0,1][ 0 , 1 ], as
$\mathbf{I}=\mathbf{I}/2^{B}$ (8)
@@
if only $\textit{Rank}([\mathbf{L},\ \mathbf{1}])\geq 2$ where $\mathbf{L}=[L_{1},L_{2},...,L_{m}]^{\text{T}}$, it would be effective to estimate the noise prior $(\sigma_{\text{s}},\ \sigma_{\text{r}})$ using a simple least square optimization algorithm.
@@
Figure 4: Framework of LoNPE network. The backbone of network is same as the one in DoTNet 26.
@@ 3.2.2 LoNPE Network
Inspired by the concept of noise representation in Huang et al 26, that the noise level could be represented as an explicit parameter and is easy to learn by an external neural network. We build a learnable CNN model to predict the noise prior $(\hat{\sigma}_{\text{s}},\hat{\sigma}_{\text{r}})\rightarrow(\sigma_{\text{s}},\sigma_{\text{r}})$ from a single sRGB noisy color image $\mathbf{I}_{\text{rgb}}$, which is formulated as
@@
where $\Phi_{\text{LoNPE}}(\cdot)$ denotes the noise prior estimation network, and called “LoNPE network”. The framework of LoNPE network is shown in Fig. 4, consisting of a backbone as the DoTNet 26 for feature extraction and two fully-connected layers for decision. Specifically, we sample only 8 random local patches from each image to estimate an accurate noise prior, instead of using $m$ patches as previously described. Due to the physical characteristics of shot noise (photon-to-electron conversion rate) and read noise (quantization range of digital signal), the output range of $\sigma_{\text{s}}$ and $\sigma_{\text{r}}$ are limited to [0,1][0,1][ 0 , 1 ].